OpenStudy (anonymous):

Let the integral I-4 (from -4.5 to -9) f(x)dx=1, the I-9 integral from -7.5 to -9 =9 and the integral from -4.5 to -6 = 9 Find integral from -6 to -7.5

OpenStudy (amistre64):

cant make sense of what you wrote

OpenStudy (anonymous):

let me try and write it out in a equation just one second thanks. :)

OpenStudy (amistre64):

k, most likely tho, they have defined .. looks like 3 areas and you should be able to just add or subtract as needed

OpenStudy (anonymous):

\[\int\limits_{-9}^{-4.5} f(x)dx=1, \int\limits_{-9}^{-7.5}f(x)dx=9, \int\limits_{-6}^{-4.5}f(x)dx=9 Find \int\limits_{-7.5}^{-6} f(x)dx\]

OpenStudy (amistre64):

from -4.5 to -9 = 1 from -7.5 to -9 =9 from -4.5 to -6 = 9 Find integral from -6 to -7.5 id add 9 to all the intervals just to get the stuff out of the negatives from 4.5 to 0 = 1 from 1.5 to 0 =9 from 4.5 to 3 = 9 Find integral from 3 to 1.5 that at least gets us out of the negative stuff

OpenStudy (amistre64):

then id do some flipping to get the limits looking right; when we flip limits, we negate the results from 0 .. 4.5 = -1 from 0 .. 1.5 = -9 from 3 .. 4.5 = -9